A projectile is projected from the origin with a velocity of 20.0 m/s at an angle of 30.0 degrees above the horizontal. The projectile lands on a roof that is 4.00 m in height. What is the time when the projectile hits the roof?

1.67 s

1.33 s

1.49 s

1.11 s

1.27 s
Answer #1

By applying kinematics law in vertical direction,

H = u*t + 0.5*a*t^2

here, H = height of building = 4.00 m

u = vertical velocity = 20.0*sin(30.0 deg)

u = 10.0 m/s

a = -9.81 m/s^2

So, 4 = 10.0*t + 0.5*(-9.81)*t^2

by solving quadratic equation,

t = 0.54 sec.and t = 1.49 sec.

therefore correct option is (C).

Please upvote.